# Engine-fed Kilonovae (Mergernovae) - II. Radiation

### v1

@@ 2.3 X-ray irradiation
As has been discussed in Paper I, the ejecta is radiation-dominated. Ignoring heating and emission when calculating the photons’ diffusion, a general equation based on energy conservation can be written as
$T^{\prime}\frac{Ds_{\rm rad}^{\prime}}{Dt^{\prime}}=-\frac{1}{\rho^{\prime}r^{2}}\frac{\partial}{\partial r}\left(r^{2}F_{\rm rad}^{\prime}\right)$ (49)
where
$s_{\rm rad}^{\prime}=\frac{4aT^{\prime 3}}{3\rho^{\prime}}$ (50)
is the specific entropy of a certain fluid element in the ejecta. $F_{\rm rad}^{\prime}$ represents the diffusive flux, which reads as
$F_{\rm rad}^{\prime}=-\frac{c}{3\rho^{\prime}\kappa}\frac{\partial u^{\prime}}{\partial r}$ (51)
where $u^{\prime}=aT^{4\prime}$ is the energy density for the radiation-dominated fluid. Based on Equation 49, the evolution of energy density can be expressed as
$\frac{Du^{\prime}}{Dt^{\prime}}=\frac{1}{r^{2}}\frac{\partial}{\partial r}\left(r^{2}\frac{c}{3\rho^{\prime}\kappa}\frac{\partial u^{\prime}}{\partial r}\right)+\frac{4u^{\prime}}{3\rho^{\prime}}\frac{D\rho^{\prime}}{Dt^{\prime}}.$ (52)
The second term on the right-hand-side of equation 52 represents the effect of $pdV$ expansion. From equation 51 one can clearly see that the diffusion of photons requires a gradient of energy density.
However, the timescale for photons to effectively diffuse across different shells depends on the optical depth in between and is evolving, which is not always comparable with the time resolution adopted to calculate the dynamics. Therefore, in this work, instead of numerically solving Equation 52, we use a simplified treatment to approximate the photon diffusion.
@@
where $l=1/(\kappa\rho^{\prime})$ represents the mean free path of photons in these shells and ${\cal N}_{s}$ is the expected number of scatterings (absorptions) for a photon to propagate through these shells. For the $n$th shell, within a time step $dt^{\prime}$, the photon can diffuse across $k_{n,f}$ shells forwards and $k_{n,b}$ shells backwards. The expected number of scattering for a photon to propagate across the shells can be estimated as
@@
and ${\cal N}_{s,n-k_{n,f}\leftrightarrow n}$ as the same, where $\delta\tau=\kappa\rho^{\prime}\delta vt$. $\kappa$ represents the opacity. $k_{n,f}$ and $k_{n,b}$ can be obtained by equaling $dt^{\prime}$ with $t_{\rm diff}$. Suppose at each time step $dt^{\prime}$, from the $(n-k_{n,b})$th to $(n+k_{n,f})$th shells, local thermal equilibrium tends to be achieved. Therefore, the luminosity of photons diffusing across the interface between the $n$th and $(n+1)$th shells is calculated as
$L_{\rm diff,n\rightarrow n+1}^{\prime}=\frac{\sum_{i=n-k_{n,b}}^{n}\delta E_{\rm int,i}^{\prime}-\sum_{i=n-k_{n,b}}^{n+k_{n,f}+1}\delta E_{\rm int,i}^{\prime}\frac{\sum_{i=n-k_{n,b}}^{n}\delta V_{i}^{\prime}}{\sum_{i=n-k_{n,b}}^{n+k_{n,f}+1}\delta V_{i}^{\prime}}}{dt^{\prime}}.$ (55)
@@
Considering all heating, radiation and photon diffusion terms introduced above, the overall evolution of the ejecta’s internal energy could be expressed as
$\frac{d(\delta E_{\rm int,n}^{\prime})}{dt^{\prime}}$ $=$ $\delta L_{\rm sh}^{\prime}+\delta L_{X,t,n}^{\prime}+\delta L_{\rm ra,n}^{\prime}$ (64)
$-\delta L_{\rm out,n}^{\prime}-\delta L_{e,n}^{\prime}-p_{n}\frac{d(\delta V_{n}^{\prime})}{dt^{\prime}}.$
The evolution of the volume of each shell is calculated with Equation 43. In the ejecta’s comoving frame, lab frame and observer’s frame, the times intervals’ relation can be written as $dt^{\prime}=dt/\Gamma={\cal D}_{N}dt_{\rm obs}$.
@@ 3 Results
The rotational kinetic energy of the newly born magnetar can be estimated as $E_{\rm rot}=\frac{1}{2}I\Omega_{0}^{2}$, where $I$ and $\Omega_{0}$ represent its momentum of inertia and initial rotational angular velocity. Adopting the fiducial values that $I=1.5\times 10^{45}~{}{\rm g~{}cm^{2}}$ and $\Omega_{0}=2\pi/P_{0}=6.3\times 10^{3}~{}{\rm rad/s}$ with $P_{0}=1~{}{\rm ms}$ as the initial spin period, $E_{\rm rot}$ should be on order of $10^{52}~{}{\rm erg}$. However, the “effective” total rotational kinetic energy of the magnetar could be smaller, either through the release of secular gravitational waves (Fan et al., 2013a; Gao et al., 2016) or due to collimation of magnetar wind energy to the jet axis (Wang et al., 2024c). In general, it is expected that $E_{\rm tot}\leq E_{\rm rot}$. Therefore, we choose several $E_{\rm tot}$ that are equal to or less than $10^{52}~{}{\rm erg}$ for our discussion.
@@
  • Case I: The spin-down luminosity is relatively large. An early blue bump in the light curve would be expected. The luminosity at a later stage is determined by total energy injection budget.
@@
Figure 1: The light curves of engine-fed kilonovae are shown in different panels, assuming different total wind energy budgets with a fixed initial spin-down luminosity of $L_{\rm sd,0}=10^{47}~{}{\rm erg/s}$. The fraction of the wind luminosity dissipated into X-rays is set to be $\eta_{X}=10^{-2}$. Across all panels, the following parameters are consistently adopted: $M_{\rm ej}=10^{-2}M_{\odot}$, $\kappa=1~{}{\rm cm^{2}~{}g^{-1}}$, $v_{\rm min}=0.05c$, and $v_{\rm max}=0.2c$. Each panel is further divided into two sub-panels, the upper and lower, which present light curves in distinct observational bands and the heating rates of different channels, respectively. The black dotted lines represent the bolometric luminosity of a typical r-process kilonova without extra energy injection. The heating channels include radioactive heating ($L_{\rm ra}$), shock heating ($L_{\rm sh}$), and X-ray irradiation ($L_{X,t}$). $L_{X}$ represents the total X-ray luminosity released from the NS wind.
@@
To gain a deeper understanding of the brightness of an engine-fed kilonova signal, we present the peak time and peak luminosity of the kilonova’s light curves in Figure 3. As can be seen, a larger total energy budget of the magnetar’s wind generally leads to a higher peak bolometric luminosity. The peak time is primarily determined by the spin-down luminosity of the post-merger magnetar, where a high spin-down luminosity tends to result in an earlier peak, because the magnetar’s wind might accelerate the ejecta and make it become transparent earlier. Varying the X-ray irradiation efficiency $\eta_{X}$ has little effect on the peak times, indicating that X-ray irradiation barely alters the ejecta dynamics. It is worth noting that a higher spin-down luminosity does not necessarily result in a higher peak luminosity or an earlier peak time of the engine-fed kilonova, particularly when the spin-down timescale of the magnetar is significantly shorter than the peak time of the light curve of the engine-fed kilonova.
@@
  • At $t\sim 0.5$ day after the merger, the peak luminosity can reach more than $10^{44}~{}{\rm erg/s}$ when $E_{\rm tot}\gtrsim 10^{51}~{}{\rm erg}$ and $L_{\rm sd,0}\gtrsim 10^{47}~{}{\rm erg/s}$. When $L_{\rm sd,0}<10^{46}~{}{\rm erg/s}$, the bolometric luminosity of the engine-fed kilonova would be close to $10^{42}~{}{\rm erg/s}$, rougly at the same level of the typical r-process kilonova.
  • At $t\sim 1$ day after the merger, the bolometric luminosity of an engine-fed kilonova also reaches a high level ($\gtrsim 10^{43}~{}{\rm erg/s}$) in the case where $L_{\rm sd,0}\lesssim 10^{46}~{}{\rm erg/s}$. The bolometric luminosity in the case where $L_{\rm sd,0}>\sim 10^{47}~{}{\rm erg/s}$ decreases to a level comparable with the typical r-process one when $\eta_{X}\lesssim 10^{-4}$, but remains relatively high when $\eta_{X}\gtrsim 10^{-2}$.
@@
Figure 4: Contours for the bolometric luminosity under different parameter sets at $0.5$ day, $1$ day and $1$ week, respectively. The left and right panels represent the cases when $\eta_{X}=10^{-2}$ and $10^{-4}$, respectively. Other parameters utilized are same as those in Figure 1.
@@ 3.2 Spectral energy distribution
With a relatively low spin-down luminosity of $L_{\rm sd,0}\sim 10^{45}~{}{\rm erg/s}$, there is almost no difference between the SED of the engine-fed kilonova and the typical r-process kilonova before they both become effectively transparent at approximately $1$ day. Subsequently, due to the continuous energy injection from the magnetar wind, the cooling process of the engine-fed kilonova is slower, resulting in an overall temperature that remains higher than a typical r-process kilonova. Consequently, in the relatively late stage, the engine-fed kilonova appears bluer the r-process one.
Figure 5: The Spectral energy distributions of (engine-fed) kilonovae at different times. The total wind energy budget is adopted to be $E_{\rm tot}=10^{52}~{}{\rm erg}$. Other parameters utilized are same as those in Figure 1.
@@
Figure 6: Similar as Figure 1 but with $L_{\rm sd,0}=10^{48}~{}{\rm erg/s}$ considered. A shock breakout signal at $0.1{\rm keV}$ is predicted.
@@ 4 Conclusion and discussion
If a long-lived NS (instead of a promptly collapsing black hole or a short-lived hyper-massive NS) is produced after a BNS merger, the NS wind would inject extra energy into the merger ejecta and power an engine-fed kilonova. We have developed a code to calculate the multi-band light curves of an engine-fed kilonova. The dynamics of the wind-ejecta interaction is based on the mechanical model of magnetized relativistic blastwaves, which has been described in detail in the first paper of this series (Ai et al., 2022). Taking into account the heating effects due to both the shock exited in the ejecta and the X-ray irradiation from somewhere behind the ejecta, we calculated the photon diffusion inside the ejecta and the thermal radiation emitted from it.
@@
  • In the case where the spin-down luminosity is relatively low, for instance, $L_{\rm sd,0}\sim 10^{45}~{}{\rm erg/s}$, the early blue bump would not be evident. Instead, the prolonged shock propagation and X-ray irradiation would lead to a substantial brightening of the kilonova in the relatively late stage. The SED of the engine-fed kilonova would resemble that of the typical r-process kilonova before approximately $1$ day, but it would be significantly bluer than the latter in the late stages, due to the continuous energy injection.

### v2

@@ 2.4 Photon diffusion
The diffusion of photons across various layers in radiation-dominated materials can be described using the heat conduction equation (Mihalas & Mihalas, 1984). Ideally, one should concurrently solve the equations governing the dynamics and photon diffusion. However, the timescale for photons to effectively diffuse between shells depends on the evolving optical depth, which is not always commensurate with the time resolution used to calculate the dynamics. The timescale for photons to diffuse across a width, characterized by an initial velocity gradient $\delta v$, can be approximately estimated as $t_{{\rm diff},\delta v}^{\prime}\sim(l/c)(\delta\tau)^{2}$ when $\delta\tau>1$, where $\delta\tau$ represents the optical depth corresponding to $\delta v$, and $l=1/(\kappa\rho^{\prime})$ is the mean free path of the photons. Given that $\delta\tau=\kappa\rho^{\prime}t\delta v$ and $\rho^{\prime}\propto v^{-\alpha}t^{-3}$, we can derive that $t_{{\rm diff},\delta v}^{\prime}\propto v^{-\alpha}t^{-1}$. It is clear that the diffusion timescale for outer layers is significantly shorter than that for inner layers, and both decrease as the merger ejecta expands. In other words, the diffusion rate for the inner layers is higher, and this rate increases as the ejecta expands. Consequently, the time step for numerically solving the heat conduction equation should be adjusted at each instant. In this study, rather than numerically solving the heat conduction equation, we employ a simplified approach to approximate photon diffusion.
@@
where ${\cal N}_{s}$ is the expected number of scatterings (absorptions) for a photon to propagate through these shells. For the $n$th shell, within a time step $\Delta t^{\prime}$, the photon can diffuse across $k_{n,f}$ shells forwards and $k_{n,b}$ shells backwards. The expected number of scattering for a photon to propagate across the shells can be estimated as
@@
and ${\cal N}_{s,n-k_{n,f}\leftrightarrow n}$ as the same, where $\delta\tau=\kappa\rho^{\prime}\delta vt$. $\kappa$ represents the opacity. $k_{n,f}$ and $k_{n,b}$ can be obtained by equaling $\Delta t^{\prime}$ with $t_{\rm diff}$. Suppose at each time step $\Delta t^{\prime}$, from the $(n-k_{n,b})$th to $(n+k_{n,f})$th shells, local thermal equilibrium tends to be achieved. Therefore, the luminosity of photons diffusing across the interface between the $n$th and $(n+1)$th shells is calculated as
$L_{\rm diff,n\rightarrow n+1}^{\prime}=\frac{\sum_{i=n-k_{n,b}}^{n}\delta E_{\rm int,i}^{\prime}-\sum_{i=n-k_{n,b}}^{n+k_{n,f}+1}\delta E_{\rm int,i}^{\prime}\frac{\sum_{i=n-k_{n,b}}^{n}\delta V_{i}^{\prime}}{\sum_{i=n-k_{n,b}}^{n+k_{n,f}+1}\delta V_{i}^{\prime}}}{\Delta t^{\prime}}.$ (51)
@@
When calculating photon diffusion, we only focus on the shocked and unshocked ejecta, as the other system components contribute little. Consequently, there exist two boundaries for the ejecta: the innermost and outermost layers. Although photons are able to diffuse across different shells of the ejecta, for the kilonova photons, the innermost and the outermost boundaries are assumed to be “walls.” In other words, we have
$\delta L^{\prime}_{\rm out,N}=-L_{\rm diff,N-1\rightarrow N}^{\prime}$ (53)
and
$\delta L^{\prime}_{\rm out,1}=L_{\rm diff,1\rightarrow 2}^{\prime}.$ (54)
The leakage of photons from the boundary is effectively addressed by incorporating the radiation term ($\delta L_{e}^{\prime}$) into Equation 2. Note that these two boundaries are “free” for X-ray photons, allowing them to pass freely through the ejecta’s boundaries.
@@ 2.5 Radiation
Considering all heating, radiation and photon diffusion terms introduced above, the overall evolution of the ejecta’s internal energy could be calculated following Equation 2. The evolution of the volume of each shell is described with Equation 43. In the ejecta’s comoving frame, lab frame and observer’s frame, the times intervals’ relation can be written as $dt^{\prime}=dt/\Gamma={\cal D}_{N}dt_{\rm obs}$.
2.6 Numerical resolution
To simulate photon diffusion and radiation processes, we have segmented the ejecta into $N$ shells, representing the spatial resolution. Fundamentally, spatial and temporal resolutions are independent parameters. However, we found that having the FS traverse one shell at each time step simplifies the calculations. Thus, in this work, the spatial and temporal resolutions are correlated.
Given the initial selection of the number of shells ($N$) for the ejecta, before the breakout of the FS, the time resolution ($\Delta t$) needs to be updated at each instant. Specifically, it is determined as $\Delta t=\Delta_{s}/(v_{f}-v_{s})$, where $\Delta_{s}=t\delta v_{s}$ represents the shell thickness that the FS is traversing in the lab frame. After the FS breaks out of the ejecta, we fix $\Delta t$ to the value it had when the FS traversed the outermost shell. It is worth noting that this fixed time resolution is defined on a logarithmic scale.
To achieve a precise light curve, adopting an appropriate time resolution is crucial. However, manually setting $\Delta t$ is challenging, as even with a fixed $N$, the time resolution varies across different parameter sets for the magnetar wind and the merger ejecta. Although we cannot formulate an analytical relation between spatial and temporal resolution, their positive correlation is undisputed. Initially, we choose $N=1000$ shells for each parameter set and follow the aforementioned model to calculate the luminosity evolution. The calculations begin at $t_{c}$ and terminate at $t_{\rm end}=10^{7}~{}{\rm s}$, with the timeline divided into roughly uniform intervals on a logarithmic scale. We set a minimum of $N_{t}=300$ time intervals as our criterion. If the results yield a lower time resolution, we increase the spatial resolution until the time resolution criteria is met. Furthermore, we test convergence by increasing $N_{t}$ to $500$, but find no significant difference in the light curve compared to when $N_{t}=300$ is used.
@@ 3 Results
The rotational kinetic energy of the newly born magnetar can be estimated as $E_{\rm rot}=\frac{1}{2}I\Omega_{0}^{2}$, where $I$ and $\Omega_{0}$ represent its momentum of inertia and initial rotational angular velocity. Adopting the fiducial values that $I=1.5\times 10^{45}~{}{\rm g~{}cm^{2}}$ and $\Omega_{0}=2\pi/P_{0}=6.3\times 10^{3}~{}{\rm rad/s}$ with $P_{0}=1~{}{\rm ms}$ as the initial spin period, $E_{\rm rot}$ should be on order of $10^{52}~{}{\rm erg}$. Due to the possible release of energy through secular gravitational waves, the total energy budget carried by the magnetar wind ($E_{\rm tot}$) could be smaller than the rotational kinetic energy ($E_{\rm rot}$). Considering the beaming effect of the magnetar wind, the “effective” total wind energy could be even smaller, denoted as $E_{\rm eff}=\eta E_{\rm tot}$. Therefore, we choose several values of $E_{\rm tot}$ that are equal to or less than $10^{52}~{}{\rm erg}$ for our discussion. Consequently, the corresponding $E_{\rm eff}$ is less than $10^{51}~{}{\rm erg}$, as $\eta=0.1$ is adopted throughout this paper.
The initial spin-down luminosity ($L_{\rm sd,0}$) and the efficiency of the wind to dissipate into X-rays ($\eta_{X}$) are not completely free parameters. In principle, they can be inferred through the observation of X-ray plateaus on the afterglow light curve of the short GRBs (Rowlinson et al., 2010, 2013; Lü et al., 2015). According to the short GRB sample with an X-ray internal plateau collected in Lü et al. (2015), the break times of the plateaus concentrate around several hundred seconds, indicating that the spin-down luminosity should be $L_{\rm sd,0}\sim{\rm a~{}few}\times 10^{49}~{}{\rm erg/s}$ if the total rotational kinetic energy is $E_{\rm rot}\sim 10^{52}~{}{\rm erg}$. The observed X-ray plateau luminosity in the Lü et al. (2015) sample range from approximately $10^{46}~{}{\rm erg/s}$ to $10^{49}~{}{\rm erg/s}$. Incorporating a beaming factor of $\eta=0.1$ (because the effective X-ray luminosity along the short GRB direction could be enhanced by a factoe of $\eta^{-1}=10$), the range for $\eta_{X}$ should be from $10^{-4}$ to ${\rm a~{}few}\times 10^{-2}$, with the majority of the sources having $\eta_{X}\sim 10^{-2}$. For $\eta_{X}\ll 1$, the dissipation of magnetic field would not influence the dynamics of the system, so that we will not modify the formulae related to the mechanical model of the blastwave. When $\eta_{X}\sim 10^{-2}$, the contribution of X-ray irradiation would be comparable with the shock heating, because the shock heating efficiency obtained from Paper I is also $\xi_{t}\sim 0.01$. In this work, we present the cases with $\eta_{X}=10^{-2}$ and $10^{-4}$ separately. Recent observations also discovered some X-ray internal plateaus sources that are not associated with short GRBs, likely due to a relatively large observing angle (as expected theoretically (Zhang, 2013)). These plateaus have an even lower inferred luminosities (Xue et al., 2019; Sun et al., 2019; Ai & Zhang, 2021b). This indicates that either a smaller $L_{\rm sd,0}$ or a smaller $\eta_{X}$ is allowed. For GW170817, if a long-lived post-merger NS is produced, the required spin-down luminosity should be even much smaller (Ai et al., 2018).
@@ 3.1 Light curve
  • Case I: The spin-down luminosity is relatively large. An early blue bump in the light curve would be expected. The luminosity at a later stage is determined by total energy injection budget and the X-ray heating efficiency $\eta_{X}$.
@@
Figure 1: The light curves of engine-fed kilonovae are shown in different panels, assuming different total wind energy budgets with a fixed initial spin-down luminosity of $L_{\rm sd,0}=10^{47}~{}{\rm erg/s}$. The fraction of the wind luminosity dissipated into X-rays is set to be $\eta_{X}=10^{-2}$ for the upper three panels and $10^{-4}$ for the lower three panels. Across all panels, the following parameters are consistently adopted: $M_{\rm ej}=10^{-2}M_{\odot}$, $\kappa=1~{}{\rm cm^{2}~{}g^{-1}}$, $v_{\rm min}=0.05c$, and $v_{\rm max}=0.2c$. Each panel is further divided into two sub-panels, the upper and lower, which present light curves in distinct observational bands (“U”,“R” and “H” in the legend represent the observational bands) and the heating rates of different channels, respectively. The solid lines represent the bolometric luminosities of the engine-fed kilonovae. The black dotted lines represent the bolometric luminosity of a typical r-process kilonova without extra energy injection. The heating channels include radioactive heating ($L_{\rm ra}$), shock heating ($L_{\rm sh}$), and X-ray irradiation ($L_{X,t}$). $L_{X}$ represents the total X-ray luminosity released from the NS wind.
@@
To gain a deeper understanding of the brightness of an engine-fed kilonova signal, we present the peak time and peak luminosity of the kilonova’s light curves in Figure 3. As can be seen, a larger total energy budget of the magnetar’s wind generally leads to a higher peak bolometric luminosity. The peak time is primarily determined by the spin-down luminosity of the post-merger magnetar, where a high spin-down luminosity tends to result in an earlier peak, because the magnetar’s wind might accelerate the ejecta and make it become transparent earlier. Varying the X-ray irradiation efficiency $\eta_{X}$ has little effect on the peak times, indicating that X-ray irradiation barely alters the ejecta dynamics. It is worth noting that a higher spin-down luminosity does not necessarily result in a higher peak luminosity of the engine-fed kilonova, particularly when the spin-down timescale of the magnetar is significantly shorter than the peak time of the light curve of the engine-fed kilonova.
@@
  • At $t\sim 0.5$ day after the merger, the peak luminosity can reach more than $10^{44}~{}{\rm erg/s}$ when $E_{\rm tot}\gtrsim 4\times 10^{51}~{}{\rm erg}$ and $L_{\rm sd,0}\gtrsim 10^{47}~{}{\rm erg/s}$. When $L_{\rm sd,0}<2\times 10^{46}~{}{\rm erg/s}$, the bolometric luminosity of the engine-fed kilonova would be close to $10^{42}~{}{\rm erg/s}$, rougly at the same level of the typical r-process kilonova.
  • At $t\sim 1$ day after the merger, the bolometric luminosity of an engine-fed kilonova also reaches a high level ($\gtrsim 10^{43}~{}{\rm erg/s}$) in the case where $L_{\rm sd,0}\lesssim 10^{46}~{}{\rm erg/s}$. The bolometric luminosity in the case where $L_{\rm sd,0}>\sim 4\times 10^{47}~{}{\rm erg/s}$ decreases to a level comparable with the typical r-process one when $\eta_{X}\lesssim 10^{-4}$, but remains relatively high when $\eta_{X}\gtrsim 10^{-2}$.
@@
Figure 4: The contours representing the bolometric luminosity under different total energy budgets and initial spin-down luminosities at 0.5 day, 1 day, and 1 week, respectively, are shown with a grey color map. The contours for the spin-down timescale ($t_{\rm sd}$) are shown with dashed lines. The left and right panels represent the cases when $\eta_{X}=10^{-2}$ and $10^{-4}$, respectively. Other parameters utilized are same as those in Figure 1.
@@ 3.2 Spectral energy distribution
With a relatively low spin-down luminosity of $L_{\rm sd,0}\sim 10^{45}~{}{\rm erg/s}$, there is almost no difference between the SED of the engine-fed kilonova and the typical r-process kilonova before they both become effectively transparent at approximately $1$ day. Subsequently, due to the continuous energy injection from the magnetar wind, the cooling process of the engine-fed kilonova is slower, resulting in an overall temperature that remains higher than a typical r-process kilonova. Consequently, in the relatively late stage, the engine-fed kilonova appears bluer than the r-process one.
Figure 5: The Spectral energy distributions of (engine-fed) kilonovae at different times. The left and right panels represent the cases when $\eta_{X}=10^{-2}$ and $10^{-4}$, respectively. The total wind energy budget is adopted to be $E_{\rm tot}=10^{52}~{}{\rm erg}$. Other parameters utilized are the same as those in Figure 1.
@@ 3.3 Shock breakout
Figure 6: Similar as Figure 1 but with $L_{\rm sd,0}=5\times 10^{48}~{}{\rm erg/s}$ considered. A shock breakout signal at $0.1{\rm keV}$ is predicted. The left and right panels represent the cases when $\eta_{X}=10^{-2}$ and $10^{-4}$, respectively.
@@ 4 Conclusion and discussion
If a long-lived NS (instead of a promptly collapsing black hole or a short-lived hyper-massive NS) is produced after a BNS merger, the NS wind would inject extra energy into the merger ejecta and power an engine-fed kilonova. We have developed a code to calculate the multi-band light curves of an engine-fed kilonova. The dynamics of the wind-ejecta interaction is based on the mechanical model of magnetized relativistic blastwaves, which has been described in detail in the first paper of this series (Ai et al., 2022). Taking into account the heating effects due to both the shock excited in the ejecta and the X-ray irradiation from somewhere behind the ejecta, we calculated the photon diffusion inside the ejecta and the thermal radiation emitted from it.
@@
  • In the case where the spin-down luminosity is relatively low, for instance, $L_{\rm sd,0}\sim 10^{45}~{}{\rm erg/s}$, the early blue bump would not be evident. Instead, the prolonged shock propagation and X-ray irradiation would lead to a substantial brightening of the kilonova in the relatively late stage. The SED of the engine-fed kilonova would resemble that of the typical r-process kilonova before approximately $1$ day, but it would be significantly bluer than the typical r-process kilonova in the late stages, due to the continuous energy injection.
@@
There is another model in the literature that predicts an early blue bump on the kilonova light curve, resulting from the ejecta being shocked by the GRB jet (Wu et al., 2022; Hamidani et al., 2024a, b), without the introduction of a post-merger magnetar. The signal powered by the GRB jet has a strong angular dependence, which is concentrated in the nearly polar direction. By contrast, our spin-down luminosity-powered, engine-fed kilonova can be observed in a wider range of observing angles.